Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability problem

Expert replies
by mrinal2100 » Tue Aug 26, 2008 3:46 am
A) Two cards are drawn at random from a wel suffled pack of cards.given than both the cards are red,what is the probability that bothe cards are:-

1)having same no
<1> 9/325 <2> 64/325 <3> 144/325 <4> 28/325


2) diff no's on them and diff suits?

<1> 72/325 <2> 81/325<3> 144/325 <4> 28/325


B)four digit no's are formed using the digits 0 to 4 without repetion.the probabillity that a no so formed is divisible by 2 is

<1> 5/8 <2> 3/8 <3> 7/9 <4>2/9


ans: for A(1)-9/325
ans-for A(2)-72/325
ans:-for B-5/8
Join the discussion
Source: — Problem Solving |

Re: probability problem

by parallel_chase » Tue Aug 26, 2008 4:37 am
mrinal2100 wrote:A) Two cards are drawn at random from a wel suffled pack of cards.given than both the cards are red,what is the probability that bothe cards are:-

1)having same no
<1> 9/325 <2> 64/325 <3> 144/325 <4> 28/325
2 suits - 26 red cards - 18 cards with same number

The question doesnt metion about that A, K, J, Q cannot be selected meaning this is not a GMAT question.

Probability of choosing first card = 18/26
Probability of choosing second card with same number = 1/25

total = 18/26 * 1/25 = 9/325

2) diff no's on them and diff suits?

<1> 72/325 <2> 81/325<3> 144/325 <4> 28/325
I think the answer choices are wrong.

probability of first card = 18/26
probability of different second card + different suit = 8/13

total=72/169
B)four digit no's are formed using the digits 0 to 4 without repetion.the probabillity that a no so formed is divisible by 2 is

<1> 5/8 <2> 3/8 <3> 7/9 <4>2/9
0,1,2,3,4

total number of digits that can be formed = 4*4*3*2=96
number of digits ending with 0 =4*3*2=24
number of digits ending with 2,4=3*3*2*2=36
(24+36)/96 = 5/8
Join the discussion